### v1

@@ Results
Additionally, our results confirm the observations made in [5] that the accuracy of results on AntiCP 2.0 (Dataset A) is higher than that on AntiCP 2.0 (Dataset B). This may be attributed to the fact that ACPs exhibit greater compositional and structural similarities to AMPs, which are the negative class in AntiCP 2.0 (Dataset B), than to random peptides, which serve as the negative class in AntiCP 2.0 (Dataset A). As a result, it is observed in our numerical experiments for model selection (see Supporting Information) that models trained on AntiCP 2.0 (Dataset A) demonstrate better performance, irrespective of the chosen model or feature encoding technique.
@@
iACP-DRLF [4] 93.00 89.60 96.40 0.86
AntiCP 2.0 [5] 92.01 92.27 91.75 0.84
iACP-FSCM [6] 88.90 87.60 90.20 0.78
AntiMF [7] 91.03 88.24 93.78 0.83
ACP-MHCNN [8] 90.00 86.50 93.30 0.80
AntiCP [35] 89.95 89.69 90.20 0.80
ACP-DL [36] 88.10 86.00 90.20 0.76
GRCI-Net [37] 87.60 87.00 88.10 0.75
ACPred [38] 85.31 87.11 83.51 0.71
ACPred-Fuse [39] 78.87 64.43 93.30 0.60
iACP [40] 77.58 78.35 76.80 0.55
PEPred-Suite [41] 57.47 40.21 74.74 0.16
@@
iACP-DRLF [4] 77.50 80.70 74.30 0.55
AntiMF [7] 75.94 65.34 86.63 0.53
AntiCP 2.0 [5] 75.43 77.46 73.41 0.51
ACP-DL [36] 74.60 74.90 74.30 0.49
GRCI-Net [37] 74.60 75.40 73.70 0.49
ACP-MHCNN [8] 74.00 73.70 74.30 0.48
ACPred-Fuse [39] 68.90 69.19 68.60 0.38
iACP [40] 55.10 77.91 32.16 0.11
PEPred-Suite [41] 53.49 33.14 73.84 0.08
ACPred [38] 53.47 85.55 21.39 0.09
AntiCP [35] 50.58 100.00 01.16 0.07
@@ Discussion
In this paper, we propose a topology-enhanced machine learning model (Top-ML) combining peptide spectral features with Magnus and natural vector-based peptide features. We demonstrate that our Top-ML model outperforms or achieves the same level of accuracy as several existing state-of-the-art machine learning models for anticancer peptide classification. Our results highlight the potential of mathematics-assisted peptide-based features to improve the performance of anticancer peptide identification.
In spite of the overall increase in predictive accuracy, there are several limitations associated with this study. Additionally, since this study focuses only on the topological featurization of peptides, it has not considered physicochemical properties such as hydrophobicity and helicity, which have been shown to also provide valuable signals in encoding a peptide. These different classes of feature encoding methods can be combined in future research.
@@
In the following experiment, we systematically tested two definitions of $f(w)$ in Equation (1) for constructing Laplacian matrices: the number of occurrences of amino acid $w$ (frequency), and its mean index position. We compared the performance of two construction methods using three types of classifiers: Random Forest (RF), Extra Trees (ETrees), and Gradient Boosting Trees (GBT). Models are evaluated on the benchmark dataset AntiCP 2.0, using peptide spectral features derived from $L_{0}$ and $L_{1}$ as inputs.
@@
Table 2: The performance of models using peptide spectral features on AntiCP 2.0 (Datasets A and B)
@@
In this experiment, detailed in Table 3, we focus on the Extra Trees classifier and evaluate its performance using the mean Magnus vector across various window sizes. A window size of 4 yields the highest accuracy on AntiCP 2.0 (Dataset A), whereas a window size of 5 achieves the highest accuracy on AntiCP 2.0 (Dataset B). We selected a window size of 5 for our final Top-ML model because it achieves the highest AUROC for both AntiCP 2.0 Datasets A and B, the highest accuracy of 76.54% on AntiCP 2.0 Dataset B, and the second-highest accuracy of 92.17% on AntiCP 2.0 Dataset A.
@@
Table 3: The performance of models using Magnus vector representation on AntiCP 2.0 (Datasets A and B)
@@
In this section, we tested three classifiers, i.e., RF, ETrees, and GBT, with the natural vector as input features. The Gradient Boosting Trees classifier provides the best accuracy of 92.53% on AntiCP 2.0 (Dataset A) (Table 4). On AntiCP 2.0 (Dataset B), Extra Trees achieves the highest accuracy of 74.78%. For both Magnus representation and natural representation, the specificity metric generally trends higher than sensitivity across different classifiers.
Table 4: Performance of models using natural vector-based features on AntiCP 2.0 (Datasets A and B)
@@
To test which terminus to use, we generate the natural vector and mean Magnus vector based on the terminus instead of the entire peptide sequence and test the performance using the Extra Trees classifier. For AntiCP 2.0 (Dataset A) (Table 5), the C15-based natural and Magnus vectors outperformed all other terminal composition encoding with an accuracy score of 92.78%. On AntiCP 2.0 (Dataset B) shown in Table 5, the N15C15 terminus composition feature provided the best results with an accuracy of 77.13%.
@@
Table 5: The performance of models using Terminal composition features on AntiCP 2.0 (Datasets A and B)
@@
Acknowledgments
@@
Code and Data Availability
The Top-ML code can be found in https://github.com/XueGong-git/TopPep-ML. The datasets for ACPs can be found in https://webs.iiitd.edu.in/raghava/anticp2/download.php.
Author contributions
K.X. designed the experiment. J.T., J.W., and X.G. conducted the experiment and analyzed the data. J.T. wrote the initial draft of the manuscript. K.X., J.W., and X.G. subsequently revised the manuscript. All authors reviewed and contributed to the final manuscript.

### v4

@@ Results
Additionally, our results confirm the observations made in 5 that the accuracy of results on AntiCP 2.0 (Dataset A) is higher than that on AntiCP 2.0 (Dataset B). This may be attributed to the fact that ACPs exhibit greater compositional and structural similarities to AMPs, which are the negative class in AntiCP 2.0 (Dataset B), than to random peptides, which serve as the negative class in AntiCP 2.0 (Dataset A). As a result, it is observed in our numerical experiments in Model Selection that models trained on AntiCP 2.0 (Dataset A) demonstrate better performance, irrespective of the chosen model or feature encoding technique.
@@
iACP-DRLF 4 93.00 89.60 96.40 0.86
AntiCP 2.0 5 92.01 92.27 91.75 0.84
iACP-FSCM 6 88.90 87.60 90.20 0.78
AntiMF 7 91.03 88.24 93.78 0.83
ACP-MHCNN 8 90.00 86.50 93.30 0.80
AntiCP 40 89.95 89.69 90.20 0.80
ACP-DL 41 88.10 86.00 90.20 0.76
GRCI-Net 42 87.60 87.00 88.10 0.75
ACPred 43 85.31 87.11 83.51 0.71
ACPred-Fuse 44 78.87 64.43 93.30 0.60
iACP 45 77.58 78.35 76.80 0.55
PEPred-Suite 46 57.47 40.21 74.74 0.16
@@
iACP-DRLF 4 77.50 80.70 74.30 0.55
AntiMF 7 75.94 65.34 86.63 0.53
AntiCP 2.0 5 75.43 77.46 73.41 0.51
ACP-DL 41 74.60 74.90 74.30 0.49
GRCI-Net 42 74.60 75.40 73.70 0.49
ACP-MHCNN 8 74.00 73.70 74.30 0.48
ACPred-Fuse 44 68.90 69.19 68.60 0.38
iACP 45 55.10 77.91 32.16 0.11
PEPred-Suite 46 53.49 33.14 73.84 0.08
ACPred 43 53.47 85.55 21.39 0.09
AntiCP 40 50.58 100.00 01.16 0.07
Model performance on mACPpred2.0
To demonstrate the generalizability of our method, we tested the model on a more recent and comprehensive ACP dataset, mACPpred 2.0 47. This dataset was compiled from 11 existing studies, excluding amino acid sequences shorter than 5 or longer than 50 residues. It also accounts for sequence similarities by removing redundancy using the CD-HIT technique with a similarity threshold of 0.85. The training dataset comprises 1,176 ACPs and 1,176 non-ACPs, while the test dataset consists of 610 ACPs and 2,760 non-ACPs. We selected the N10C10 terminus to construct terminal composition features for this dataset, and the performance of our model compared to other benchmark models is presented in Table 2 below. Our tree-based model achieves similar accuracy when compared to other more complex deep learning models such as mACPpred 2.0 47 and MLACP 2.0 48. Both of these models use complex deep learning framework that integrate probabilistic features computed from various baseline models. or example, mACPpred2.0 49 uses a stacked deep learning approach, combining multiple base learners trained on different feature types with a 1D CNN that integrates the predictions of base models into the final output. Its features also include 16 pre-trained NLP-based embeddings, leveraging external datasets. In contrast, our model employs vectors directly derived from amino acid sequences and uses a tree-based classifier, offering better computational efficiency and greater interpretability, as demonstrated by the feature importance analysis in Figure 3.
Table 2: Performance comparison of Top-ML with existing state-of-the-art machine learning models on mACPpred 2.0.
Model Acc Sen Spc MCC
mACPpred 2.0 47 85.7 80.7 86.8 59.7
MLACP 2.0 48 83.8 79.2 84.9 55.7
AntiCP 2.0 (Alternate) 5 83.6 60.5 88.7 47.2
Top-ML 82.5 74.4 84.3 51.4
iDACP (Sp of 90%) 50 82.5 20.3 96.3 25.5
iDACP (Sp of 80%) 50 82.4 35.6 92.7 33.0
iDACP (Sp of 70%) 50 81.8 48.4 89.2 38.0
PreTP-2L 51 81.8 2.1 99.4 6.5
iDACP (Sp of 60%) 50 81.7 61.6 86.1 44.0
CancerGram 52 81.5 6.6 98.1 11.0
iAMP-RAAC 53 81.4 6.6 97.9 10.2
mACPpred 49 80.7 75.4 81.9 48.9
iDACP (Sp of 50%) 50 80.5 68.0 83.3 45.1
AMPfun 54 76.0 76.9 75.8 42.9
ACPred-BMF (Alternate) 55 75.9 44.1 82.9 25.1
ACPred 43 75.1 51.0 80.4 27.8
TriNet 56 74.7 43.3 81.7 22.9
ACP-MHCNN (740) 8 71.1 35.7 78.9 13.2
AntiCP 2.0 (Main) 5 71.0 43.3 77.2 17.8
ACP-MHCNN (500/164) 8 61.6 16.7 71.6 -10.2
ACPred-BMF (Main) 55 60.0 67.2 58.4 19.8
iAMPCN 57 24.1 62.5 15.6 -21.3
In Figure 3a, we present the feature importance calculated from the Top-ML model using the Gini importance criterion, which measures the total reduction in Gini impurity contributed by each feature. To understand how these feature values vary between non-ACPs and ACPs, we identified the three highest peaks in Figure 3a and plotted the distribution of the feature values for each class in Figure 3b. The largest spike corresponds to a feature extracted from the natural vector constructed from the N10C10 terminus vector, representing the mean position of the amino acid “E” in the N10C10 terminus. On average, the amino acid “E”, or glutamic acid, appears later in non-ACPs than in ACPs. The second-highest spike corresponds to a spectrum from the Hodge Laplacian $\tilde{\mathbf{L}}_{0}$, where non-ACPs exhibit a larger spectrum than ACPs. The spectrum of $\tilde{\mathbf{L}}_{0}$ is associated with clustering patterns 58, where smaller values indicate that amino acids from the same clusters are more likely to appear adjacent to each other. he plot suggests that ACPs exhibit a stronger clustering pattern than non-ACPs. The rightmost figure displays a feature corresponding to the third-highest peak, derived from the mean Magnus vector, which represents the occurrence of glutamic acid. The plot suggests that glutamic acid is more frequent in non-ACPs than in ACPs. his observation is consistent with the analyses presented in 50, 40. This could be explained by the fact that ACPs tend to have high hydrophobicity and a positive net charge to effectively interact with negatively charged cancer cells 50, whereas glutamic acid is negatively charged and hydrophilic.
Figure 3: Feature importance of the Top-ML model, averaged over 100 iterations of training using the mACPpred 2.0 training set. Feature importance is computed using the Gini importance criterion. We identified the three highest peaks in the top plot, highlighted with markers, and plotted their value distributions for each class in the bottom plot.
@@ Discussion
In this paper, we propose a topology-enhanced machine learning model (Top-ML) that combines peptide spectral features with Magnus and natural vector-based peptide features. We demonstrate that our Top-ML model based on an Extra Trees classifier outperforms or achieves similar level of accuracy as several existing state-of-the-art deep learning models for ACPs classification on benchmark datasets AntiCP 2.0 and mACPpred 2.0. Our results highlight the potential of mathematics-assisted peptide-based features to improve the performance of ACPs identification. Furthermore, our model provides improved interpretability, demonstrated through feature importance analysis on the mACPpred 2.0 dataset.
Since this study focuses only on the topological featurization of peptides, it has not considered physicochemical properties such as hydrophobicity and helicity, which have been shown to also provide valuable signals in encoding a peptide. ncorporating the 3D structures, for instance through spectral features computed using the simplicial complex representation, could be an interesting direction for future research. Our framework focus on developing expressive feature extraction methods for sequential data. Therefore, while our experiments focused on anticancer peptides due to data availability and their biomedical significance, the framework is inherently applicable to various sequence types, including amino acid (proteins and peptides) and nucleotide (DNA and RNA) sequences. Future research could investigate its generalizability and effectiveness across other datasets and various prediction or regression tasks. For instance, most existing peptide datasets consist of natural peptides. Expanding the framework to include modified peptides could offer valuable insights into their potential in drug application.
@@
In the following experiment, we systematically tested two definitions of $f(w)$ in Equation (1) for constructing Laplacian matrices: the number of occurrences of amino acid $w$ (frequency), and its mean index position. e compared the 5-fold cross-validation performance on the training data of two construction methods using three types of classifiers : Random Forest (RF), Extra Trees (ETrees), and Gradient Boosting Trees (GBT). Models are evaluated on the benchmark dataset AntiCP 2.0, using peptide spectral features derived from $L_{0}$ and $L_{1}$ as inputs.
@@
Table 3: The 5-fold cross-validation performance using peptide spectral features on AntiCP 2.0 (Datasets A and B)
@@
In this 5-fold cross-validation experiment on the training data, detailed in Table 4, we focus on the Extra Trees classifier and evaluate its performance using the mean Magnus vector across various window sizes. A window size of 4 yields the highest accuracy on AntiCP 2.0 (Dataset A), whereas a window size of 5 achieves the highest accuracy on AntiCP 2.0 (Dataset B). We selected a window size of 5 for our final Top-ML model because it achieves the highest AUROC for both AntiCP 2.0 Datasets A and B, the highest accuracy of 76.54% on AntiCP 2.0 Dataset B, and the second-highest accuracy of 92.17% on AntiCP 2.0 Dataset A.
@@
Table 4: The 5-fold cross-validation performance using Magnus vector representation on AntiCP 2.0 (Datasets A and B)
@@
In this section, we tested three classifiers, i.e., RF, ETrees, and GBT, using the natural vector as input features ith 5-fold cross-validation on the training data. The Gradient Boosting Trees classifier provides the best accuracy of 92.53% on AntiCP 2.0 (Dataset A) (Table 5). On AntiCP 2.0 (Dataset B), Extra Trees achieves the highest accuracy of 74.78%. For both Magnus representation and natural representation, the specificity metric generally trends higher than sensitivity across different classifiers.
Table 5: The 5-fold cross-validation performance using natural vector-based features on AntiCP 2.0 (Datasets A and B)
@@
To test which terminus to use, we generate the natural vector and mean Magnus vector based on the terminus instead of the entire peptide sequence and test the performance using the Extra Trees classifier ith 5-fold cross validation on the training data. For AntiCP 2.0 (Dataset A) (Table 6), the C15-based natural and Magnus vectors outperformed all other terminal composition encoding with an accuracy score of 92.78%. On AntiCP 2.0 (Dataset B) shown in Table 6, the N15C15 terminus composition feature provided the best results with an accuracy of 77.13%.
@@
Table 6: The 5-fold cross-validation performance using Terminal composition features on AntiCP 2.0 (Datasets A and B)
@@
Data and Software Availability
The Top-ML code can be found in https://github.com/XueGong-git/TopPep-ML. The datasets for antiCP 2.0 can be found in https://webs.iiitd.edu.in/raghava/anticp2/download.php. The mACPpred 2.0 datasets are available from https://github.com/nhattruongpham/mACPpred2.
Author Contributions
K.X. designed the experiments. J. T., J.W. and X.G. conducted the experiments and analyzed the results. J.T. drafted the initial manuscript. J.W. and X.G. revised the manuscript. All authors reviewed and contributed to the final manuscript.
Conflict of Interest
The authors declare no competing financial interests.
{acknowledgement}